# CDFlow: Building Invertible Layers with Circulant and Diagonal Matrices

### v1

@@ 1 Introduction
Deep generative models have gained significant attention due to their potential applications across various domains, including natural language processing (Dong et al., 2024), image restoration and generation (Yao et al., 2023), robot learning (Chao et al., 2024), and autonomous driving (Guan et al., 2024). Among these models, normalizing flows hold a distinctive position because of their ability to provide accurate likelihood estimation and efficient sampling.
@@
To address this challenge, NICE (Dinh et al., 2014) introduced additive coupling layers, which were later extended by RealNVP (Dinh et al., 2016) to affine coupling layers. These coupling layers have since become a fundamental component of most flow-based models. However, the design of these non-linear layers, which only transform a subset of variables, requires coordination with linear layers to effectively capture dependencies across arbitrary dimensions. As a result, recent efforts have focused on developing specialized linear layer designs, ranging from fixed channel-disrupting operations (Dinh et al., 2014, 2016) to $1\times 1$ convolutions (Kingma and Dhariwal, 2018) and even $d\times d$ convolutions (Hoogeboom et al., 2019) with learnable parameters. However, dense convolutions can result in significant memory overhead as a result of the large weight matrices required for high-dimensional inputs. One approach to address this is the construction of reversible butterfly layers (Meng et al., 2022) using butterfly matrices (Dao et al., 2019), which help maintain the model’s expressiveness while reducing the number of parameters.
Structured matrices and Fourier transform acceleration have been demonstrated in numerous studies (Cheng et al., 2015; Ding et al., 2017; Karami et al., 2019; Jensen et al., 2021) to be highly effective in reducing the number of parameters and computational complexity of neural networks. Research by (Huhtanen and Perämäki, 2015) demonstrates that any matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$ can be represented as the alternating product of a cyclic matrix and a diagonal matrix, requiring no more than 2n12n-1 factors. This factorization structure has demonstrated versatility and effectiveness in domains such as deep neural networks  (Moczulski et al., 2015; Araujo et al., 2019) and parameter-efficient fine-tuning  (Ding et al., 2025). However, these studies have not focused on model reversibility.
Inspired by previous work, we propose a novel normalizing flow model called Circulant-Diagonal Flow (CDFlow), which is built on invertible linear layers constructed using circulant and diagonal matrices. This construction preserves the expressive power of the model while significantly reducing storage demands, as only the diagonal and circulant elements need to be stored instead of the full $n\times n$ weight matrix. In addition, the distinct properties of circulant and diagonal matrices enable the use of the Fast Fourier Transform (FFT) to enhance computation efficiency (Ding et al., 2025). This facilitates faster calculations for both the determinant of Jacobian matrices and the inversion of weight matrices. Overall, the main contributions of our work can be summarized as following:
@@ 2.1 Flow-based Models
The coupling layer (Dinh et al., 2014, 2016) is a simple and effective invertible transformation, where the Jacobi matrix is a triangular matrix. It divides the input into two parts $x_{a}$ and $x_{b}$, then transforms $x_{b}$ using a scaling parameter $s_{\theta}$ and a translation parameter $b_{\theta}$ that can be a nonlinear function of $x_{a}$, while $z_{a}$ is directly equal to $x_{a}$, and the final output concatenates the two parts together:
@@
Coupling layers remain a crucial component in most flow-based models (Kingma and Dhariwal, 2018; Hoogeboom et al., 2019; Meng et al., 2022). However, since coupling layers only transform a subset of the variables, they must be combined with linear layers to facilitate information fusion across channels. As a result, recent research has increasingly focused on the design of linear layers to enhance model performance.
@@
Glow (Kingma and Dhariwal, 2018) pioneered the use of $1\times 1$ convolution with learnable parameters as a linear layer in flow-based models. This method is more effective at capturing correlations between variables compared to simpler channel shuffling operations (Dinh et al., 2014, 2016). Additionally, Glow introduces a weight design based on LU decomposition, reducing the Jacobian determinant computation from $\mathcal{O}(n^{3})$ to $\mathcal{O}(n)$.
Similarly, (Hoogeboom et al., 2019) extended convolution to $d\times d$ and proposed periodic convolution, leveraging the Fast Fourier Transform (FFT) to accelerate the convolution process. However, the periodic convolution method involves generalized determinant computation, leading to a complexity of $\mathcal{O}(n^{3})$. In contrast, our proposed method not only accelerates the convolution process using FFT, but also reduces the determinant computation complexity to $\mathcal{O}(n)$.
Despite numerous works exploiting special matrix properties to enhance model expressiveness, most focus on improving the efficiency of the Jacobian determinant computation, with limited optimization of the matrix inversion process during the sampling phase of the flow model. In these methods, including the aforementioned works (Kingma and Dhariwal, 2018; Hoogeboom et al., 2019; Meng et al., 2022), the computational complexity for the inverse sampling operation is typically $\mathcal{O}(n^{2})$. The Woodbury Transformation (Lu and Huang, 2020) achieves a theoretical complexity of $\mathcal{O}(dn)$, where $d$ is the dimensionality of smaller matrices (e.g., 8 or 16). Although this is better in theory than the $\mathcal{O}(n\log n)$ complexity of our method, in practice Woodbury requires multiple input transformations, which makes its actual efficiency inferior to ours.
@@ 2.2 Weight Matrix Construction and Fast Multiplication
A general matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$ can be factorized into the alternating product of circulant and diagonal matrices, with the total number of matrices not exceeding 2n12n-1 (Huhtanen and Perämäki, 2015). Specifically, it can be written as:
@@
where $\mathbf{D}_{2j-1}$ and $\mathbf{C}_{2j}$ are diagonal and circulant matrix for $j=1,2,\dots,n$. This decomposition can theoretically approximate any dense matrix, and the approximation error can be controlled by adjusting the number of factors. (Ding et al., 2025) have successfully applied such representation to fine-tuning large language models.
@@ 3.2 Matrix Inverse
$1\times 1$ Conv (Kingma and Dhariwal, 2018) $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$
$1\times 1$ Conv (LU) (Kingma and Dhariwal, 2018) $\mathcal{O}(n)$ $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$
Emerging (Hoogeboom et al., 2019) $\mathcal{O}(n)$ $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$
Periodic (Hoogeboom et al., 2019) $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$
Woodbury (Lu and Huang, 2020) $\mathcal{O}(nd)$ $\mathcal{O}(nd)$ $\mathcal{O}(d(n+h\cdot w))$
ME-Woodbury (Lu and Huang, 2020) $\mathcal{O}(nd)$ $\mathcal{O}(nd)$ $\mathcal{O}(d(n+h+w))$
ButterflyFlow (Meng et al., 2022) $\mathcal{O}(n)$ $\mathcal{O}(n^{2})$ $\mathcal{O}(nL)$
@@ 3.3 Generative Modeling with CDFlow
Figure 1: Architecture of the CDFlow Model. The left portion shows a detailed schematic of the proposed CD-Convolution layer, which is constructed using circulant and diagonal matrices. The middle portion illustrates the internal structure of a flow module, where the input sequentially passes through an Actnorm layer(Kingma and Dhariwal, 2018), the CD Convolution layer, and a coupling layer(Dinh et al., 2016). The right portion presents the overall CDFlow model architecture, in which the flow module is repeated K times and integrated into a multi-scale framework(Dinh et al., 2016) to effectively fuse channel-wise information.
@@
The ActNorm layer (Kingma and Dhariwal, 2018) is a type of normalization technique that uses affine transformations to perform class normalization. Unlike traditional methods like BatchNorm (Ioffe and Szegedy, 2015), which rely on the statistics of the entire batch, the ActNorm layer initializes its parameters based on the statistics of the first batch. Once initialized, these parameters are treated as learnable variables. This initialization method is advantageous for small batch sizes, as it avoids the performance degradation that BatchNorm often encounters when dealing with small batches. ActNorm helps in stabilizing the training of normalizing flows and is especially useful for flow-based models that require flexible scaling and translation invariance.
@@
The coupling layer (Dinh et al., 2014, 2016) is a core component of normalizing flow models. The coupling layer splits the input into two parts: one part is passed through a neural network to compute a transformation, while the other part is left unchanged. This operation enables the model to learn complex transformations while maintaining a simple and computationally efficient structure. The coupling layer is combined with the reversible convolutional layer to form the flow module, allowing the CDFlow model to capture complex distributions and generate high-quality samples.
@@
CDFlow incorporates a multi-scale architecture (Dinh et al., 2016), which improves the model’s ability to capture complex and high-dimensional data distributions. Within each block, the split and squeeze operations play an important role:
@@ 4.1 Experimental Settings and Model Parameters
To compare our method with previous approaches, we evaluate its performance on the CIFAR-10 (Krizhevsky, 2009) and ImageNet  (Deng et al., 2009) datasets, using bits per dimension (BPD) as the evaluation metric. BPD is a commonly used measure for assessing the quality of generative models, indicating the number of bits required to encode each dimension of the data. Metrics such as IS and FID are not used in this work, as they do not generalize well across different datasets and fail to account for overfitting. For the baseline, we select several models related to CDFlow, including MAF (Papamakarios et al., 2017), Real NVP (Dinh et al., 2016), Glow (Kingma and Dhariwal, 2018), Emerging (Hoogeboom et al., 2019), i-ResNet (Behrmann et al., 2019), Woodbury (Lu and Huang, 2020) and ButterflyFlow (Meng et al., 2022).
@@
During training, we apply spectral normalization (Miyato et al., 2018) to the convolutional layers to ensure stability. Thanks to the structure of circulant and diagonal matrices, the maximum singular value of each weight matrix can be efficiently computed and used to rescale the weights. In addition, we employ a channel-aware learning rate scaling for the CD-Convolution parameters, which prevents abrupt updates to the structured matrices. Empirically, this combination further stabilizes the training of CDFlow.
@@ 4.2 Density Estimation on Standard Images
To enable direct comparison with prior work (Meng et al., 2022), we evaluate our method on the CIFAR-10 (Krizhevsky, 2009) and ImageNet-32×32 (Deng et al., 2009) datasets. For both datasets, training is conducted with unified dequantization and standard data augmentation techniques.
@@ Results.
MAF (Papamakarios et al., 2017) 105.0M 4.31
Real NVP (Dinh et al., 2016) 6.4M 3.49 46.2M 4.28 6.38M 2.11
Glow (Kingma and Dhariwal, 2018) 44.2M $3.36\pm 0.002$ 66.2M 4.09 6.25M 2.02
Emerging (Hoogeboom et al., 2019) 46.6M $3.34\pm 0.002$ 44.0M 4.09 6.68M 1.98
Residual Flows (Chen et al., 2019) 25.2M 3.28 47.1M 4.01 5.58M 3.60
i-ResNet (Behrmann et al., 2019) 25.0M $3.32\pm 0.006$ 29.7M 4.05 6.88M 2.04
Woodbury (Lu and Huang, 2020) 45.3M $3.42\pm 0.002$ 45.3M 4.09 6.68M 2.01
i-DenseNet (Perugachi-Diaz et al., 2021) 24.9M 3.25 47.0M 3.98 6.07M 4.06
ButterflyFlow (Meng et al., 2022) 44.4M $3.33\pm 0.003$ 44.4M 4.09 6.39M 1.95
@@ 4.3 Density Estimation on Structured Images
To demonstrate the capability of CDFlow in modeling periodic structures, we apply it to galaxy images (Ackermann et al., 2018), conducting an experiment on this real-world dataset.
@@
When the boundary pixels of an image are relatively uniform or connected, the data can be considered periodic. A typical example of periodic data is galaxy images, which often feature scattered light sources with predominantly dark boundaries. The Galaxy datasets (Ackermann et al., 2018) is a small classification dataset that includes 3,000 merged galaxies and 5,000 non-merged galaxies in the training and test sets. We train CDFlow on non-merged galaxy images and compare its performance to other models under the same experimental setup.
@@ 4.4 Runtime
To demonstrate the effectiveness of our method in reducing computational complexity, we separately evaluated the runtime of the forward operation, log-determinant (logdet) computation, and inverse operation. We compared our method with standard $1\times 1$ convolution, emerging and periodic convolutions (Hoogeboom et al., 2019), the Woodbury transformation and ME-Woodbury with a fixed latent dimension of $d=16$ (Lu and Huang, 2020), and the butterfly layer with level 10 (Meng et al., 2022). For a fair comparison, all methods were implemented in PyTorch, and all experiments were conducted on an NVIDIA A800 GPU.

### v3

@@ 1 Introduction
Deep generative models have gained significant attention due to their potential applications across various domains, including natural language processing (dong2024dreamllm), image restoration and generation (yao2023local), robot learning (chao2024maximum), and autonomous driving (10522953). Among these models, normalizing flows hold a distinctive position because of their ability to provide accurate likelihood estimation and efficient sampling.
@@
To address this challenge, NICE (dinh2014) introduced additive coupling layers, which were later extended by RealNVP (dinh2016) to affine coupling layers. These coupling layers have since become a fundamental component of most flow-based models. However, the design of these non-linear layers, which only transform a subset of variables, requires coordination with linear layers to effectively capture dependencies across arbitrary dimensions. As a result, recent efforts have focused on developing specialized linear layer designs, ranging from fixed channel-disrupting operations (dinh2014; dinh2016) to $1\times 1$ convolutions (kingma2018) and even $d\times d$ convolutions (hoogeboom2019emerging) with learnable parameters. However, dense convolutions can result in significant memory overhead as a result of the large weight matrices required for high-dimensional inputs. One approach to address this is the construction of reversible butterfly layers (meng2022butterflyflow) using butterfly matrices (Dao2019LearningFA), which help maintain the model’s expressiveness while reducing the number of parameters.
Structured matrices and Fourier transform acceleration have been demonstrated in numerous studies (cheng2015exploration; ding2017circnn; karami2019invertible; jensen2021scalable) to be highly effective in reducing the number of parameters and computational complexity of neural networks. Research by (Huhtanen2015FactoringMI) demonstrates that any matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$ can be represented as the alternating product of a cyclic matrix and a diagonal matrix, requiring no more than 2n12n-1 factors. This factorization structure has demonstrated versatility and effectiveness in domains such as deep neural networks  (moczulski2015acdc; araujo2019understanding) and parameter-efficient fine-tuning  (ding2025parameter). However, these studies have not focused on model reversibility.
Inspired by previous work, we propose a novel normalizing flow model called Circulant-Diagonal Flow (CDFlow), which is built on invertible linear layers constructed using circulant and diagonal matrices. This construction preserves the expressive power of the model while significantly reducing storage demands, as only the diagonal and circulant elements need to be stored instead of the full $n\times n$ weight matrix. In addition, the distinct properties of circulant and diagonal matrices enable the use of the Fast Fourier Transform (FFT) to enhance computation efficiency (ding2025parameter). This facilitates faster calculations for both the determinant of Jacobian matrices and the inversion of weight matrices. Overall, the main contributions of our work can be summarized as following:
@@ 2.1 Flow-based Models
The coupling layer (dinh2014; dinh2016) is a simple and effective invertible transformation, where the Jacobi matrix is a triangular matrix. It divides the input into two parts $x_{a}$ and $x_{b}$, then transforms $x_{b}$ using a scaling parameter $s_{\theta}$ and a translation parameter $b_{\theta}$ that can be a nonlinear function of $x_{a}$, while $z_{a}$ is directly equal to $x_{a}$, and the final output concatenates the two parts together:
@@
Coupling layers remain a crucial component in most flow-based models (kingma2018; hoogeboom2019emerging; meng2022butterflyflow). However, since coupling layers only transform a subset of the variables, they must be combined with linear layers to facilitate information fusion across channels. As a result, recent research has increasingly focused on the design of linear layers to enhance model performance.
@@
Glow (kingma2018) pioneered the use of $1\times 1$ convolution with learnable parameters as a linear layer in flow-based models. This method is more effective at capturing correlations between variables compared to simpler channel shuffling operations (dinh2014; dinh2016). Additionally, Glow introduces a weight design based on LU decomposition, reducing the Jacobian determinant computation from $\mathcal{O}(n^{3})$ to $\mathcal{O}(n)$.
Similarly, (hoogeboom2019emerging) extended convolution to $d\times d$ and proposed periodic convolution, leveraging the Fast Fourier Transform (FFT) to accelerate the convolution process. However, the periodic convolution method involves generalized determinant computation, leading to a complexity of $\mathcal{O}(n^{3})$. In contrast, our proposed method not only accelerates the convolution process using FFT, but also reduces the determinant computation complexity to $\mathcal{O}(n)$.
Despite numerous works exploiting special matrix properties to enhance model expressiveness, most focus on improving the efficiency of the Jacobian determinant computation, with limited optimization of the matrix inversion process during the sampling phase of the flow model. In these methods, including the aforementioned works (kingma2018; hoogeboom2019emerging; meng2022butterflyflow), the computational complexity for the inverse sampling operation is typically $\mathcal{O}(n^{2})$. The Woodbury Transformation (Lu2020woodbury) achieves a theoretical complexity of $\mathcal{O}(dn)$, where $d$ is the dimensionality of smaller matrices (e.g., 8 or 16). Although this is better in theory than the $\mathcal{O}(n\log n)$ complexity of our method, in practice Woodbury requires multiple input transformations, which makes its actual efficiency inferior to ours.
@@ 2.2 Weight Matrix Construction and Fast Multiplication
A general matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$ can be factorized into the alternating product of circulant and diagonal matrices, with the total number of matrices not exceeding 2n12n-1 (Huhtanen2015FactoringMI). Specifically, it can be written as:
@@
where $\mathbf{D}_{2j-1}$ and $\mathbf{C}_{2j}$ are diagonal and circulant matrix for $j=1,2,\dots,n$. This decomposition can theoretically approximate any dense matrix, and the approximation error can be controlled by adjusting the number of factors. (ding2025parameter) have successfully applied such representation to fine-tuning large language models.
@@ 3.2 Matrix Inverse
$1\times 1$ Conv (kingma2018) $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$
$1\times 1$ Conv (LU) (kingma2018) $\mathcal{O}(n)$ $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$
Emerging (hoogeboom2019emerging) $\mathcal{O}(n)$ $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$
Periodic (hoogeboom2019emerging) $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$
Woodbury (Lu2020woodbury) $\mathcal{O}(nd)$ $\mathcal{O}(nd)$ $\mathcal{O}(d(n+h\cdot w))$
ME-Woodbury (Lu2020woodbury) $\mathcal{O}(nd)$ $\mathcal{O}(nd)$ $\mathcal{O}(d(n+h+w))$
ButterflyFlow (meng2022butterflyflow) $\mathcal{O}(n)$ $\mathcal{O}(n^{2})$ $\mathcal{O}(nL)$
@@ 3.3 Generative Modeling with CDFlow
Figure 1: Architecture of the CDFlow Model. The left portion shows a detailed schematic of the proposed CD-Convolution layer, which is constructed using circulant and diagonal matrices. The middle portion illustrates the internal structure of a flow module, where the input sequentially passes through an Actnorm layer(kingma2018), the CD Convolution layer, and a coupling layer(dinh2016). The right portion presents the overall CDFlow model architecture, in which the flow module is repeated K times and integrated into a multi-scale framework(dinh2016) to effectively fuse channel-wise information.
@@
The ActNorm layer (kingma2018) is a type of normalization technique that uses affine transformations to perform class normalization. Unlike traditional methods like BatchNorm (Ioffe2015BatchNA), which rely on the statistics of the entire batch, the ActNorm layer initializes its parameters based on the statistics of the first batch. Once initialized, these parameters are treated as learnable variables. This initialization method is advantageous for small batch sizes, as it avoids the performance degradation that BatchNorm often encounters when dealing with small batches. ActNorm helps in stabilizing the training of normalizing flows and is especially useful for flow-based models that require flexible scaling and translation invariance.
@@
The coupling layer (dinh2014; dinh2016) is a core component of normalizing flow models. The coupling layer splits the input into two parts: one part is passed through a neural network to compute a transformation, while the other part is left unchanged. This operation enables the model to learn complex transformations while maintaining a simple and computationally efficient structure. The coupling layer is combined with the reversible convolutional layer to form the flow module, allowing the CDFlow model to capture complex distributions and generate high-quality samples.
@@
CDFlow incorporates a multi-scale architecture (dinh2016), which improves the model’s ability to capture complex and high-dimensional data distributions. Within each block, the split and squeeze operations play an important role:
@@ 4.1 Experimental Settings and Model Parameters
To compare our method with previous approaches, we evaluate its performance on the CIFAR-10 (Krizhevsky2009LearningML) and ImageNet  (Deng2009image) datasets, using bits per dimension (BPD) as the evaluation metric. BPD is a commonly used measure for assessing the quality of generative models, indicating the number of bits required to encode each dimension of the data. Metrics such as IS and FID are not used in this work, as they do not generalize well across different datasets and fail to account for overfitting. For the baseline, we select several models related to CDFlow, including MAF (Papamakarios2017MAF), Real NVP (dinh2016), Glow (kingma2018), Emerging (hoogeboom2019emerging), i-ResNet (behrmann2019IRN), Woodbury (Lu2020woodbury) and ButterflyFlow (meng2022butterflyflow).
@@
During training, we apply spectral normalization (miyato2018spectral) to the convolutional layers to ensure stability. Thanks to the structure of circulant and diagonal matrices, the maximum singular value of each weight matrix can be efficiently computed and used to rescale the weights. In addition, we employ a channel-aware learning rate scaling for the CD-Convolution parameters, which prevents abrupt updates to the structured matrices. Empirically, this combination further stabilizes the training of CDFlow.
@@ 4.2 Density Estimation on Standard Images
To enable direct comparison with prior work (meng2022butterflyflow), we evaluate our method on the CIFAR-10 (Krizhevsky2009LearningML) and ImageNet-32×32 (Deng2009image) datasets. For both datasets, training is conducted with unified dequantization and standard data augmentation techniques.
@@ Results.
MAF (Papamakarios2017MAF) 105.0M 4.31
Real NVP (dinh2016) 6.4M 3.49 46.2M 4.28 6.38M 2.11
Glow (kingma2018) 44.2M $3.36\pm 0.002$ 66.2M 4.09 6.25M 2.02
Emerging (hoogeboom2019emerging) 46.6M $3.34\pm 0.002$ 44.0M 4.09 6.68M 1.98
Residual Flows (chen2019residualflows) 25.2M 3.28 47.1M 4.01 5.58M 3.60
i-ResNet (behrmann2019IRN) 25.0M $3.32\pm 0.006$ 29.7M 4.05 6.88M 2.04
Woodbury (Lu2020woodbury) 45.3M $3.42\pm 0.002$ 45.3M 4.09 6.68M 2.01
i-DenseNet (Perugachi2021) 24.9M 3.25 47.0M 3.98 6.07M 4.06
ButterflyFlow (meng2022butterflyflow) 44.4M $3.33\pm 0.003$ 44.4M 4.09 6.39M 1.95
@@ 4.3 Density Estimation on Structured Images
To demonstrate the capability of CDFlow in modeling periodic structures, we apply it to galaxy images (ackermann2018using), conducting an experiment on this real-world dataset.
@@
When the boundary pixels of an image are relatively uniform or connected, the data can be considered periodic. A typical example of periodic data is galaxy images, which often feature scattered light sources with predominantly dark boundaries. The Galaxy datasets (ackermann2018using) is a small classification dataset that includes 3,000 merged galaxies and 5,000 non-merged galaxies in the training and test sets. We train CDFlow on non-merged galaxy images and compare its performance to other models under the same experimental setup.
@@ 4.4 Runtime
To demonstrate the effectiveness of our method in reducing computational complexity, we separately evaluated the runtime of the forward operation, log-determinant (logdet) computation, and inverse operation. We compared our method with standard $1\times 1$ convolution, emerging and periodic convolutions (hoogeboom2019emerging), the Woodbury transformation and ME-Woodbury with a fixed latent dimension of $d=16$ (Lu2020woodbury), and the butterfly layer with level 10 (meng2022butterflyflow). For a fair comparison, all methods were implemented in PyTorch, and all experiments were conducted on an NVIDIA A800 GPU.